1. What is the reduced mass of an H35 Cl molecule? Give your answer in both kg and amu.
Answer #1

Reduced mass = √(m​​​​​​1m2/(m1+m2))

m1 (H) = 1.0078 m2 (35 Cl) = 34.9688

Reduced mass = √(1.0078×34.9688)/(1.0078+34.9688)

= 0.9796 amu

1 amu = 1.661×10-27 kg

Reduced mass = 0.9796×1.661×10-27 Kg

= 1.627×10-27 Kg

please give feedback!